# Spherical Orbital Dynamics and Relativistic Precession in Kerr-MOG Spacetime

### v1

@@ I Introduction
Building on this framework, a number of recent studies have investigated various aspects of black holes in MOG theory. In particular, the enhanced gravitational coupling has been shown to influence the formation and growth of supermassive black holes ZhoolidehHaghighi:2023kgs . Observable signatures of Kerr-MOG black holes have been extensively analyzed, including shadow morphology and light deflection Wang:2018prk ; Moffat:2019uxp ; Kuang:2022ojj , with further generalizations to Kerr-MOG-(anti-)de Sitter (AdS) spacetimes Liu:2024lbi . Higher-dimensional extensions of MOG black holes have revealed distinct geometric and thermodynamic properties Cai:2020igv , and polarized synchrotron ring images have been studied as possible observational signatures Qin:2022kaf . Alongside these efforts, ongoing advances in gravitational wave astronomy provide new opportunities to test MOG thoery in the strong-field regime.
@@
Recent Very Long Baseline Interferometry (VLBI) observations of M87* have identified periodic oscillations in the direction of its relativistic jet, consistent with Lense-Thirring precession caused by a misaligned accretion disk Cui:2023uyb . These results support earlier predictions of frame dragging around rotating black holes Lense:1918zz ; Mashhoon:1984fj , and have sparked renewed interest in relativistic precession phenomena Chakraborty:2012wv ; Chakraborty:2013naa ; Wu:2023wld ; Pradhan:2024jav ; Meng:2024gcg . In this context, analyzing spin-induced and geometric precessions in modified gravity may help identify potential deviations from GR.
@@
In this work, we conduct a detailed analysis of spherical orbits in Kerr-MOG spacetime, with particular emphasis on how the MOG parameter $\alpha$ and orbital tilt angle $\zeta$ affect the conserved quantities and relativistic precession frequencies. We examine three precessions relevant to strong-field dynamics: nodal precession arising from frame dragging, periastron precession induced by spacetime curvature, and the spin precession of a test gyroscope along the orbit. While previous studies Pradhan:2020nno ; Pradhan:2024jav have investigated how generalized and Lense-Thirring precession frequencies can distinguish between non-extremal black holes, extremal configurations, and naked singularities in Kerr-MOG spacetime, our focus lies in quantifying how these precessional effects are modulated by $\alpha$ and $\zeta$ across a broad parameter space. This analysis aims to improve our understanding of test particle dynamics in MOG thoery and to provide a theoretical basis for identifying potential observational signatures in gravitational wave astronomy and black hole phenomenology.
The paper is organized as follows. In Sec. II, we review the MOG theory and Kerr-MOG black hole solution. Sec. III is devoted to analyzing spherical orbits and their dynamical features. In Sec. IV, we compute nodal and periastron precession frequencies and investigate their dependence on orbital parameters and MOG corrections. In Sec. V, we examine the Lense-Thirring precession of a test gyroscope. We conclude in Sec. VI with a summary and discussion of future directions. Throughout this work, we adopt geometrized units with $G_{\rm N}=c=1$.
@@ III.1 Geodesic Structure and Conserved Quantities
In Kerr-MOG spacetime, the enhanced gravitational coupling $G=G_{\rm N}(1+\alpha)$ modifies the effective potentials $\mathcal{R}(r)$ and $\Theta(\theta)$ via the $\alpha$-dependent term, thereby shifting the stability conditions for spherical orbits. These modifications, along with potential vector field contributions, lead to deviations in the Carter constant relative to the Kerr case. As a result, $\mathcal{K}$ serves as a sensitive probe of orbital geometry, capturing the full three-dimensional structure of motion and encoding the dynamical signatures of MOG thoery.
@@ III.2 Spherical Orbits: Dynamical Features and MOG Effects
While $\alpha$ does not appear explicitly in these expressions, it indirectly affects $\mathcal{K}$ through its influence on $E$ and $L$.
@@
Comparing $\alpha=0.1$ (top row) and $\alpha=1$ (bottom row) in Fig. 1, we find that increasing $\alpha$ shifts the ISSO, energy, angular momentum, and the Carter constant to larger radii. This trend persists across different tilt angles. For instance, for $\zeta=\frac{\pi}{4}$ prograde motion, the ISSO shifts from $r=4.05M$ ($\alpha=0.1$) to $7.98M$ ($\alpha=1$), with $E$ increasing from 0.9075 to 0.9155. The enhanced gravity shifts stable spherical orbits to larger radii by providing greater centripetal force, and requires higher particle energy to maintain equilibrium. Consequently, both the ISSO radius and the associated energy increase with $\alpha$. At a significant distance of $r=50M$, larger $\alpha$ lowers the energy ($E=0.9806$ vs. $0.9892$) but increases both angular momentum ($L=7.40M$ vs. $5.39M$) and Carter constant ($\mathcal{K}=54.82M^{2}$ vs. $29.09M^{2}$) compared to the near-GR case.

### v2

@@ I Introduction
Building on this framework, a number of recent studies have investigated various aspects of black holes in MOG theory. In particular, the enhanced gravitational coupling has been shown to influence the formation and growth of supermassive black holes ZhoolidehHaghighi:2023kgs . Observable signatures of Kerr-MOG black holes have been extensively analyzed, including shadow morphology and light deflection Wang:2018prk ; Moffat:2019uxp ; Kuang:2022ojj , with further generalizations to Kerr-MOG-(anti-)de Sitter (AdS) spacetimes Liu:2024lbi . Higher-dimensional extensions of MOG black holes have revealed distinct geometric and thermodynamic properties Cai:2020igv , and polarized synchrotron ring images have been studied as possible observational signatures Qin:2022kaf . Moreover, the dynamics and emission mechanisms of relativistic jets in Kerr-MOG spacetime have been explored through particle motion simulations and modified Blandford–Znajek magnetospheres, revealing deviations from GR and providing prospects for strong-field tests LopezArmengol:2016nyi ; Camilloni:2023wyn . Complementary to these developments, recent work in the Kerr-Newman framework has yielded exact analytic solutions for null and timelike geodesics—including spherical, inspiraling, and plunging orbits—using elliptic integrals and Jacobian elliptic functions, offering valuable insight into the role of black hole charge and spin in shaping relativistic trajectories Wang:2022ouq ; Ko:2023igf . Alongside these efforts, ongoing advances in gravitational wave astronomy provide new opportunities to test MOG theory in the strong-field regime.
@@
Recent Very Long Baseline Interferometry (VLBI) observations of M87* have identified periodic oscillations in the direction of its relativistic jet, consistent with Lense-Thirring precession caused by a misaligned accretion disk Cui:2023uyb . These results support earlier predictions of frame dragging around rotating black holes Lense:1918zz ; Mashhoon:1984fj , and have sparked renewed interest in relativistic precession phenomena Chakraborty:2012wv ; Chakraborty:2013naa ; Wu:2023wld ; Pradhan:2024jav ; Meng:2024gcg . In this context, prior studies have explored how orbital motion around rotating black holes encodes information about the final spin of black hole mergers, offering useful guidance for modeling relativistic dynamics Jai-akson:2017ldo . Analyzing spin-induced and geometric precessions in modified gravity may help identify potential deviations from GR.
@@
In this work, we conduct a detailed analysis of spherical orbits in Kerr-MOG spacetime, with particular emphasis on how the MOG parameter $\alpha$ and orbital tilt angle $\zeta$ affect the conserved quantities and relativistic precession frequencies. We examine three precessions relevant to strong-field dynamics: nodal precession arising from frame dragging, periastron precession induced by spacetime curvature, and the spin precession of a test gyroscope along the orbit. While previous studies Pradhan:2020nno ; Pradhan:2024jav have investigated how generalized and Lense-Thirring precession frequencies can distinguish between non-extremal black holes, extremal configurations, and naked singularities in Kerr-MOG spacetime, our focus lies in quantifying how these precessional effects are modulated by $\alpha$ and $\zeta$ across a broad parameter space. This analysis aims to improve our understanding of test particle dynamics in MOG theory and to provide a theoretical basis for identifying potential observational signatures in gravitational wave astronomy and black hole phenomenology.
This paper is structured as follows. In Sec. II, we review the MOG theory and Kerr-MOG black hole solution. Sec. III is devoted to analyzing spherical orbits and their dynamical features. In Sec. IV, we compute nodal and periastron precession frequencies and investigate their dependence on orbital parameters and MOG corrections. In Sec. V, we examine the Lense-Thirring precession of a test gyroscope. We conclude in Sec. VI with a summary and discussion of future directions. Throughout this work, we adopt geometrized units with $G_{\rm N}=c=1$.
@@ III.1 Geodesic Structure and Conserved Quantities
In Kerr-MOG spacetime, the enhanced gravitational coupling $G=G_{\rm N}(1+\alpha)$ modifies the effective potentials $\mathcal{R}(r)$ and $\Theta(\theta)$ via the $\alpha$-dependent term, thereby shifting the stability conditions for spherical orbits. These modifications, along with potential vector field contributions, lead to deviations in the Carter constant relative to the Kerr case. As a result, $\mathcal{K}$ serves as a sensitive probe of orbital geometry, capturing the full three-dimensional structure of motion and encoding the dynamical signatures of MOG theory.
@@ III.2 Spherical Orbits: Dynamical Features and MOG Effects
A detailed analytical derivation of the angular momentum and Carter constant for photon orbits in Kerr-MOG spacetime can be found in Ref. BenAchour:2025uzp , offering a complementary perspective to our treatment of massive particles. While $\alpha$ does not appear explicitly in these expressions, it indirectly affects $\mathcal{K}$ through its influence on $E$ and $L$.
@@
Comparing $\alpha=0.1$ (top row) and $\alpha=1$ (bottom row) in Fig. 1, we find that increasing $\alpha$ shifts the ISSO, energy, angular momentum, and the Carter constant to larger radii. This trend persists across different tilt angles. For instance, for $\zeta=\frac{\pi}{4}$ prograde motion, the ISSO shifts from $r=4.05M$ ($\alpha=0.1$) to $7.98M$ ($\alpha=1$), with $E$ increasing from 0.9075 to 0.9155. The enhanced gravity shifts stable spherical orbits to larger radii by providing greater centripetal force and requires higher particle energy to maintain equilibrium. Consequently, both the ISSO radius and the associated energy increase with $\alpha$. At a significant distance of $r=50M$, larger $\alpha$ lowers the energy ($E=0.9806$ vs. $0.9892$) but increases both angular momentum ($L=7.40M$ vs. $5.39M$) and Carter constant ($\mathcal{K}=54.82M^{2}$ vs. $29.09M^{2}$) compared to the near-GR case.